# FragmentRetro: A Quadratic Retrosynthetic Method Based on Fragmentation Algorithms

### v1

@@ 4.3 Case studies
To demonstrate the practical utility of FragmentRetro, we evaluate drugs and natural products, as shown in Fig. 2 (and Section C). These targets are selected from the Higherlev study [13] and represent different levels of retrosynthetic difficulty: Narlaprevir, which can be solved without a higher-level strategy; Martinellic Acid, which requires a higher-level strategy to be solved; and Lennoxamine, which is not solvable by either approach. For Narlaprevir, we apply BRICS fragmentation rules, while for Martinellic Acid and Lennoxamine, we use r-BRICS since these compounds contain fused rings that BRICS would otherwise leave intact. FragmentRetro generates hundreds of solutions for each of these compounds. Fig. 2 presents a representative solution for each compound, selected to minimize the number of fragment combinations.
@@ 5 Conclusion
FragmentRetro provides a scalable framework for fragment-based retrosynthesis, achieving quadratic complexity through stock-aware fragment exploration and pruning. While FragmentRetro does not explicitly return a DAG, it performs the same core task as other retrosynthetic methods: identifying chemically valid decompositions of a target molecule into known precursors. DAG construction is a post-processing step that depends on the fragment connectivity and can be layered atop our approach. Therefore, comparing the search complexity of FragmentRetro to tree search and DMS methods is both appropriate and informative. Unlike DMS and tree-search methods that do not rely on template-based single-step models, FragmentRetro remains compatible with atom mapping tools, making it a flexible approach for retrosynthetic planning. Future work may focus on integrating cost-aware metrics and experimental validation to further refine its applicability. Additionally, mapping solutions and BBs to fully elaborated synthetic routes (form DAG) could enable direct route quality comparisons across different methods, potentially using top-$k$ accuracy as an evaluation metric.
@@ A.1 Tree search
Here, we adopt the notation introduced in the DESP paper [16] and begin by focusing on template-based tree-search retrosynthetic approaches. These approaches are compatible with any single-step retrosynthesis machine learning model, although certain models may not provide atom mappings for reactions. Let $\mathcal{M}$ denote the set of all molecules, $\mathcal{R}$ the set of all reactions, and $\mathcal{T}$ the set of all reaction templates. A reaction $R_{i}\in\mathcal{R}$ is represented as a tuple $(r_{i},p_{i},t_{i})$, where $r_{i}\subset\mathcal{M}$ is the set of reactants, $p_{i}\in\mathcal{M}$ is the product, and $t_{i}\in\mathcal{T}$ is the corresponding retro template. A retro template $t$ is a function $t:\mathcal{M}\rightarrow 2^{\mathcal{M}}$ that maps a product molecule $p_{i}$ to a set of potential reactants $r_{i}$, such that for any given reaction $R_{i}$, $r_{i}\in t_{i}(p_{i})$.
@@ A.2 DirectMultiStep
DMS methods [14] approach the retrosynthesis problem differently from tree-search methods. Instead of representing synthetic routes as sets of reactions, DMS represents an entire synthetic route as a single string in a recursive dictionary format. ML models are trained to output these route strings given a target molecule $p^{*}$ as input. Valid routes that follow the same two constraints can similarly be represented as a set of reactions $S=\{R_{1},\ldots,R_{n}\}$ that forms a DAG. The $R_{i}$ here is a tuple $(r_{i},p_{i})$ without a template.
@@
The neural network in DMS is trained to generate such recursive representations, predicting both the products and reactants for each step of the synthesis as a next token prediction task. This eliminates the need for explicit tree traversal and instead focuses on learning the structure of valid experimental routes [15] directly.
@@ B.3 FragmentRetro complexity analysis
    Substructure search: Searching for each fragment combination in the compound stock set involves substructure search, which is typically implemented in cheminformatics libraries like RDKit. The complexity of RDKit’s substructure search depends on the matching algorithm. Subgraph isomorphism has exponential complexity in the worst case. However, the use of algorithms like VF2 and Ullman algorithms can significantly bring the practical scenario for fragment molecular graphs to $\mathcal{O}(|\mathcal{B}|\cdot h^{3})$ [26]. This estimate assumes that the fragment sizes are $\mathcal{O}(h)$, though most fragments are considerably smaller. Besides, the use of SMARTS patterns and fingerprint screening significantly reduces this complexity in practice [28, 27].
@@
    Fingerprint screening: To further optimize substructure search, fingerprint screening is employed as a preprocessing step to rapidly filter out BBs that cannot contain the query substructure. Pattern fingerprints encode the presence or absence of predefined molecular fragments into bit vectors, ensuring no false negatives (where "negative" means no match). This screening step reduces the number of candidate molecules from the compound stock set before the more computationally expensive substructure matching. The number of candidates is reduced by an approximate factor of $k^{h}$, where $1<k<=2$ depends on fingerprint specificity. The optimal case of $k=2$ occurs when each heavy atom feature is hashed into the pattern fingerprint, and the fingerprint dimension is large enough to avoid double hashing. In this case, each feature eliminates half of the BBs on average, leading to $k=2$. Given these constraints, since $h$ has a natural upper bound for organic molecules and $k^{h}$ grows exponentially, $h^{3}/k^{h}$ rapidly diminishes to a near-constant value in practice, allowing us to approximate it as a constant $C_{fp}$. Intuitively, larger fragments require fewer substructure matches since they contain more 1’s in their pattern fingerprint bit vectors.
@@ Appendix C Additional case studies
Figure 3: Additional evaluation of target compounds using FragmentRetro. This figure adopts the format and compound set from Extended Data Fig. 5 of the Higher-Level Retrosynthesis paper [13] to facilitate direct comparison. The smallest box contains targets solved by the “Original” tree-search algorithm (as defined in [13]); the medium box includes additional compounds solved by the Higher-Level strategy. Compounds outside both boxes were not solved by either method. Compounds successfully solved by BRICS + FragmentRetro and r-BRICS + FragmentRetro are marked with green (BRICS+FR) and blue (r-BRICS+FR) tags, respectively.

### v2

@@ 4.2 Case studies
To demonstrate the practical utility of FragmentRetro, we evaluate drugs and natural products, as shown in Fig. 2 (and Section C). These targets are selected from the Higherlev study [higherlev_2025] and represent different levels of retrosynthetic difficulty: Narlaprevir, which can be solved without a higher-level strategy; Martinellic Acid, which requires a higher-level strategy to be solved; and Lennoxamine, which is not solvable by either approach. For Narlaprevir, we apply BRICS fragmentation rules, while for Martinellic Acid and Lennoxamine, we use r-BRICS since these compounds contain fused rings that BRICS would otherwise leave intact. FragmentRetro generates hundreds of solutions for each of these compounds. Fig. 2 presents a representative solution for each compound, selected to minimize the number of fragment combinations.
@@ 5 Conclusion
FragmentRetro provides a scalable framework for fragment-based retrosynthesis, achieving quadratic complexity through stock-aware fragment exploration and pruning. While FragmentRetro does not explicitly return a DAG, it performs the same core task as other retrosynthetic methods: identifying chemically valid decompositions of a target molecule into known precursors. Its outputs provide sets of fragment combinations that capture aspects of both single-step (individual disconnections) and multistep (multiple simultaneous combinations) approaches, although the method does not define the exact sequence of reactions or BBs. DAG construction is a post-processing step that depends on the fragment connectivity and can be layered atop our approach. Therefore, comparing the search complexity of FragmentRetro to tree search and DMS methods is both appropriate and informative. Unlike DMS and tree-search methods that do not rely on template-based single-step models, FragmentRetro remains compatible with atom mapping tools, making it a flexible approach for retrosynthetic planning. However, FragmentRetro is not a route-level synthesis planning tool in its current form. Mapping solutions and BBs to fully elaborated synthetic routes (forming a DAG) would be required to enable direct route quality comparisons across different methods, potentially using top-$k$ accuracy as an evaluation metric. Future work may focus on integrating cost-aware metrics and experimental validation to further refine its applicability.
@@ A.1 Tree search
Here, we adopt the notation introduced in the DESP paper [desp_2024] and begin by focusing on template-based tree-search retrosynthetic approaches. These approaches are compatible with any single-step retrosynthesis machine learning model, although certain models may not provide atom mappings for reactions. Let $\mathcal{M}$ denote the set of all molecules, $\mathcal{R}$ the set of all reactions, and $\mathcal{T}$ the set of all reaction templates. A reaction $R_{i}\in\mathcal{R}$ is represented as a tuple $(r_{i},p_{i},t_{i})$, where $r_{i}\subset\mathcal{M}$ is the set of reactants, $p_{i}\in\mathcal{M}$ is the product, and $t_{i}\in\mathcal{T}$ is the corresponding retro template. A retro template $t$ is a function $t:\mathcal{M}\rightarrow 2^{\mathcal{M}}$ that maps a product molecule $p_{i}$ to a set of potential reactants $r_{i}$, such that for any given reaction $R_{i}$, $r_{i}\in t_{i}(p_{i})$.
@@ A.2 DirectMultiStep
DMS methods [dms_2025] approach the retrosynthesis problem differently from tree-search methods. Instead of representing synthetic routes as sets of reactions, DMS represents an entire synthetic route as a single string in a recursive dictionary format. ML models are trained to output these route strings given a target molecule $p^{*}$ as input. Valid routes that follow the same two constraints can similarly be represented as a set of reactions $S=\{R_{1},\ldots,R_{n}\}$ that forms a DAG. The $R_{i}$ here is a tuple $(r_{i},p_{i})$ without a template.
@@
The neural network in DMS is trained to generate such recursive representations, predicting both the products and reactants for each step of the synthesis as a next token prediction task. This eliminates the need for explicit tree traversal and instead focuses on learning the structure of valid experimental routes [paroutes] directly.
@@ B.3 FragmentRetro complexity analysis
    Substructure search: Searching for each fragment combination in the compound stock set involves substructure search, which is typically implemented in cheminformatics libraries like RDKit. The complexity of RDKit’s substructure search depends on the matching algorithm. Subgraph isomorphism has exponential complexity in the worst case. However, the use of algorithms like VF2 and Ullman algorithms can significantly bring the practical scenario for fragment molecular graphs to $\mathcal{O}(|\mathcal{B}|\cdot h^{3})$ [substruct_algo_2004]. This estimate assumes that the fragment sizes are $\mathcal{O}(h)$, though most fragments are considerably smaller. Besides, the use of SMARTS patterns and fingerprint screening significantly reduces this complexity in practice [substruct_bench_2012, rdkit_2006].
@@
    Fingerprint screening: To further optimize substructure search, fingerprint screening is employed as a preprocessing step to rapidly filter out BBs that cannot contain the query substructure. Pattern fingerprints encode the presence or absence of predefined molecular fragments into bit vectors, ensuring no false negatives (where "negative" means no match). This screening step reduces the number of candidate molecules from the compound stock set before the more computationally expensive substructure matching. The number of candidates is reduced by an approximate factor of $k^{h}$ ($k>1$). This occurs when each heavy atom feature is hashed into the pattern fingerprint, and the fingerprint dimension is large enough to avoid double hashing. In practical scenarios, $k$ is often around 2 , when each feature eliminates roughly half of the BBs on average. Given these constraints, since $h$ has a natural upper bound for organic molecules and $k^{h}$ grows exponentially, $h^{3}/k^{h}$ rapidly diminishes to a near-constant value in practice, allowing us to approximate it as a constant $C_{fp}$. Intuitively, larger fragments require fewer substructure matches since they contain more 1’s in their pattern fingerprint bit vectors.
@@ Appendix C Additional case studies
Figure 3: Additional evaluation of target compounds using FragmentRetro. This figure adopts the format and compound set from Extended Data Fig. 5 of the Higher-Level Retrosynthesis paper [higherlev_2025] to facilitate direct comparison. The smallest box contains targets solved by the “Original” tree-search algorithm (as defined in [higherlev_2025]); the medium box includes additional compounds solved by the Higher-Level strategy. Compounds outside both boxes were not solved by either method. Compounds successfully solved by BRICS + FragmentRetro and r-BRICS + FragmentRetro are marked with green (BRICS+FR) and blue (r-BRICS+FR) tags, respectively.